### v1

@@ 3 Model Construction
We thus have a sample of stars in the halo, which can be considered as complete in the northern hemisphere at $r_{\rm gc}\lesssim 50$ kpc after correction of the selection function. That is, we have the distribution function (DF) of halo stars in 6D phase space $f(\boldsymbol{x},\boldsymbol{v})$ numerically known from the observations. We take the assumption that the smooth halo is stationary; that the DF of stars integrated in the correct potential matches that directly observed instantaneously. And under this assumption, the DF constructed from the northern hemisphere should be representative of the whole system.
@@
This method, in principle, is similar to the traditional Schwarzschild orbit-superposition method, but we have the true DF (orbit library and orbit weights) known from the observations. The only assumption we take in this model is that the system is stationary, which is the minimum assumption for all dynamical models. And we will confirm the feasibility of this assumption for the smooth halo with the good match between model and data, as we will show in Section 3.4.
@@ 3.1 Gravitational potential
in which $R_{\rm gc}=\sqrt{x_{\rm gc}^{2}+y_{\rm gc}^{2}}$. Our data in the halo do not have strong constraints on the bulge and disk. We fix the density distribution of the bulge and disks by the true values from the simulation (Grand et al., 2017). We have $M_{*,\rm bulge}=3.2\times 10^{10}M_{\odot}$, scale radius $a_{\rm bulge}=1.71$ kpc, and sérsic index $n=1.44$, $M_{*,\rm disk}=5.6\times 10^{10}M_{\odot}$, scale radius $a_{\rm disk}=4.9$ kpc, scale height $b_{\rm disk}=0.25$ kpc for Auriga 23. The disk and bulge mass still have some uncertainty for the MW. We try models with disk mass $20\%$ higher or lower for Auriga 23 and find no difference in our results. The true values parameterize the bulge and disk of Auriga 5 and Auriga 12 in Grand et al. (2017) are adopted.
We adopt a flexible triaxial generalized NFW model for the DM halo, allowing free orientations of the DM halo, and variable 3D shape as a function of radius generally following Vasiliev et al. (2021):
@@
The coordinate $\boldsymbol{X}\equiv\{X,Y,Z\}$ is related to the galactocentric Cartesian coordinate $\boldsymbol{x}_{\rm gc}\equiv\{x_{\rm gc},y_{\rm gc},z_{\rm gc}\}$ by a rotation matrix parametrized by three Euler angles $\alpha_{q}$, $\beta_{q}$, $\gamma_{q}$ as shown in Section 2.2. In principle, all three angles can be left free. Considering the degeneracy between DM orientation and 3D shape, we fix $\beta_{q}=0$. The combination of $\alpha_{q}$ and $\gamma_{q}$ determines the DM halo orientation in the $XY$ plane; we fix $\gamma_{q}=0$ and allow $\alpha_{q}$ to be free to test the ability of constraining it. For the default models, we also fix $\alpha_{q}=0$.
@@
For the radial profile, we fixed $\alpha=1$, $\beta=3$, chose outer cutoff radius $r_{\rm cut}=500$ kpc, and cutoff strength $\xi=5$. We start with models for all three mock galaxies with fixed DM orientation ($\alpha_{q}=0$, $\beta_{q}=0$ and $\gamma_{q}=0$) and constant $p_{\rm DM}$ and $q_{\rm DM}$. The five parameters in the DM halo: the scale density $\rho_{0}$ (or the mass $M_{\rm halo}$), scale radius $r_{s}$, inner density slope $\gamma$, $p_{\rm DM}$, and $q_{\rm DM}$ are left as free parameters. For Auriga 12, we further construct a model with different orientation angle $\alpha_{q}$ and a model allowing $p(r)$ and $q(r)$ to vary as a function of radius.
@@ 3.2 Orbit integration
Orbit integration is performed using the publicly released AGAMA package (Vasiliev, 2019) [^1] https://github.com/GalacticDynamics-Oxford/Agama. We have about 20,000 stars in the cleaned sample of the smooth halo, with their positions and velocities measured from observations, and the weight of each star obtained from the correction of selection. We start with the position and velocity of the stars from observations and integrate 10 orbital periods for each star. We withdraw 1000 particles from each orbit with equal time steps, and each particle is given the weight as the star that initializes the orbit. We superpose particles drawn from the orbits together, thus obtaining an orbit-superposed model, which numerically represents the DF of the stellar system.
@@ 3.3 Comparing model and data
The DF of the orbit-superposed model, $f_{\rm model}(\boldsymbol{x},\boldsymbol{v}|\rm potential)$, is based on data but incorporates the gravitational potential. If the system is in dynamical equilibrium, the DF of the model with the accurate gravitational potential ought to statistically be the same as the stars that initialized it $f_{\rm data}(\boldsymbol{x},\boldsymbol{v})$. We detail particular components of the DF: the spatial density distribution and velocity distributions at different positions. We evaluate the goodness of the model by comparing these against the data.
We calculate the stellar density distribution in the $R_{\rm gc}$-$z_{\rm gc}$ plane, by dividing it into $25\times 25$ bins within 50 kpc. The density in each bin is calculated by $N/(2\pi RdRdz)$ in units of $N/{\rm kpc}^{3}$. We calculate the density directly for both the observed data $\rho_{\rm data}$ and the model $\rho_{\rm model}$. The uncertainty of the data $d\rho_{\rm data}$ is assessed by bootstrapping the position of stars with their uncertainties. The density distributions of the data and model are shown in Fig. 4; we calculate the $\chi^{2}_{\rm den}$ by directly comparing the data and model:
@@
To analyze velocity distributions, we categorize model particles into $N_{\rm bin}=7\times 6$ bins in $r$ and $\theta$. The divisions are $r=[5,8,12,16,22,30,40,50]$ kpc and $\theta$=[0,10,20,30,40,55,90]. We then construct the velocity distributions ($v_{r}$, $v_{\phi}$, $v_{\theta}$) within each bin, as illustrated in Fig. 3. For each velocity component within bin $j$, we represent its distribution by a histogram using $M$ intervals, labeled as ($v^{k}_{m,j}$, $h^{k}_{m,j}$), where $k$ corresponds to the velocities $v_{r}$, $v_{\phi}$, and $v_{\theta}$.
@@
The collective likelihood of all $N_{j}$ stars in this bin is given by $L_{j}^{k}=\sum_{i=1}^{N_{\rm j}}\log(P^{k}_{i,j})$, and the overall likelihood across all the bins is represented by $L^{k}=\sum_{j=7}^{N_{\rm bin}}L_{j}^{k}$. Due to a deficit of stars at $r<4$ kpc, which contribute to the distribution function (DF) of the inner bins, the DF for the innermost bins of our model remains incomplete. Therefore, we omit the first 6 bins within the range r=[5,8] in our likelihood calculations. The likelihood across the three velocity components is merged to derive the constraints from the velocity distributions:
@@ 3.4 Best-fitting models
We investigate the parameter space of $\rho_{0}$, $r_{s}$, $\gamma$, $p_{\rm DM}$, and $q_{\rm DM}$ within the gravitational potential. A parameter grid is established with intervals of 0.1, 5 kpc, 0.02, 0.04, and 0.04 for $\rho_{0}$, $r_{s}$, $\gamma$, $p_{\rm DM}$, and $q_{\rm DM}$ respectively. An iterative approach is employed to search for optimal models. The process begins with an initial model and, after developing initial models, iterative refinement follows. During each iteration, we identify the optimal models using the criterion $\chi^{2}-\min(\chi^{2})<\chi^{2}_{s}$, where $\chi^{2}_{s}=100$. We then generate new models by walking two steps in every direction of the parameter grid from each optimal model. This approach guides the search towards lower $\chi^{2}$ within the parameter grid and halts when the model with the minimum $\chi^{2}$ is identified. The relatively high value of $\chi^{2}_{s}$ is selected to ensure that all models within $1\sigma$ confidence level are considered before the end of the iteration. Ultimately, we determine the best-fitting models by achieving the minimum $\chi^{2}_{\rm tot}$.
@@
We illustrate how the model works using Auriga 23 as an example. In Fig. 4, it is shown that the best-fitting model accurately reproduces the stellar density across the $R_{\rm gc}-z_{\rm gc}$ plane. Furthermore, it imposes strict constraints on the 3D shape of the underlying dark DM halo, since variations in the DM shape can cause the constructed stellar halo to appear overly flat or excessively round. As shown in Fig. 5, the best-fitting model also well reproduces the velocity distributions in all three components: $v_{r}$, $v_{\phi}$, and $v_{\theta}$. These velocity distributions, particularly $v_{r}$, provide strong constraints on the radial density distribution of the DM halo, as differing radial profiles of the DM halo can lead to notable discrepancies between observational data and the model. Our best-fitting model succeeds in reproducing both the density and velocity distributions at various locations across the $R_{\rm gc}-z_{\rm gc}$ plane, thereby validating our assumption of a ’stationary system’ with the high agreement between the model and data.
@@
In Fig. 6, we illustrate the parameter space explored for Auriga 23. The colorful points denote models within a $3\sigma$ confidence level, colored by the normalized $\Delta\chi^{2}_{\rm tot}=(\chi^{2}_{\rm tot}-\min(\chi^{2}_{\rm tot}))/\Delta\chi^{2}_{1\sigma}$, where $\Delta\chi^{2}_{1\sigma}$ is the $1\sigma$ confidence level determined by bootstrapping 222 We take the gravitational potential of the best-fitting model, and perturb the position and velocity of observed stars with their uncertainty for 100 trials. In each trial, we recompute the model, and evaluate the resulting $\chi^{2}_{\rm tot}$ (including $\chi^{2}_{\rm den}$ and $\chi^{2}_{\rm vdis}$). The standard deviation of these 100 $\chi^{2}$ values is taken as $1\sigma$ confidence level of the model $\Delta\chi^{2}_{1\sigma}$.. As shown in Fig. 6, the combination of density distribution and velocity distributions strongly constrains the DM 3D shape $p_{\rm DM}$ and $q_{\rm DM}$. Due to the still limited data coverage, there are still strong degeneracies between the three parameters ($\rho_{0}$, $r_{s}$, and $\gamma$) determining the radial profile. Nevertheless, the radial density profile within the data coverage is well constrained, as will be detailed in the next section.
@@ 4.1 Recovery of DM density profile
We demonstrate how effectively our models recover the DM radial density profile and the enclosed DM mass in Fig. 7. Despite significant degeneracy among $\rho_{0}$, $r_{s}$, and $\gamma$, our models accurately capture the radial density profile within the observational range of 4-50 kpc. Uncertainty is larger at $r<8$ kpc due to the absence of data.
The enclosed DM mass profile is also well recovered, with a relative uncertainty of about $10\%$. It is worth noting that the DM mass we obtained in the outer regions is $5\%-10\%$ higher than the true values but aligns well with the enclosed mass of DM and gas combined. The CircumGalactic Medium (CGM) in the halo of the simulation mirrors the DM’s density distribution. The mass of CGM is compensated by an overestimate of DM mass without a separate gas contribution in our model of the gravitational potential.
@@ 4.2.1 Models with constant $q_{\rm DM}$ and $p_{\rm DM}$
Initially, we developed models with constant $q_{\rm DM}$ and $p_{\rm DM}$ for all three galaxies. As depicted in Fig. 8, our model successfully recovered both $p_{\rm DM}$ and $q_{\rm DM}$ for Auriga 23 and Auriga 5, with relatively small uncertainties. For Auriga 12, the model captured the outer DM halo configuration, and there is a significant degeneracy between $p_{\rm DM}$ and $q_{\rm DM}$ when the halo is vertically aligned, characterized by $p_{\rm DM}<q_{\rm DM}$.
The stellar density distribution in the $R_{\rm gc}-z_{\rm gc}$ plane, especially the flattening of the stellar halo, is crucial for constraining the DM shape. We measure the stellar halo flattening, $q_{\rm star}$, across various radii using iso-density contours and compared these from the data and several models in Fig. 9. The stellar halo of Auriga 23 is notably flat, becoming increasingly so from the center to the outskirts, and our best-fitting model successfully mimicked $q_{\rm star}$ at all radii. The stellar halo of Auriga 5 appears rounder, with a tendency to become rounder moving outward; our model approximates the radial variation $q_{\rm star}(r)$, albeit with moderate inner region discrepancies. The stellar halo of Auriga 12 is the most spherical, notably rounding from inner to outer areas, with $q_{\rm star}(r)$ challenging to replicate by a model with a constant DM shape. The failure of reproducing the tendency of $q_{\rm star}(r)$ suggests a variation of DM shapes from the inner to outer regions.

### v2

@@ 3 Model Construction
We thus have a sample of stars in the halo, which can be considered complete in the northern hemisphere at $r_{\rm gc}\lesssim 50$ kpc after correction of the selection function. That is, we have the distribution function (DF) of halo stars in 6D phase space $f(\boldsymbol{x},\boldsymbol{v})$ numerically known from the observations. We take the assumption that the smooth halo is stationary; that the DF of stars integrated in the correct potential matches that directly observed instantaneously. And under this assumption, the DF constructed from the northern hemisphere should be representative of the whole system.
@@
This method, in principle, is similar to the traditional Schwarzschild orbit-superposition method, but we have the true DF (orbit library and orbit weights) known from the observations. The only assumption we make in this model is that the system is stationary, which is the minimum assumption for all dynamical models. And we will confirm the feasibility of this assumption for the smooth halo with a good match between the model and the data, as we will show in Section 3.4.
@@ 3.1 Gravitational potential
in which $R_{\rm gc}=\sqrt{x_{\rm gc}^{2}+y_{\rm gc}^{2}}$. Our data in the halo do not have strong constraints on the bulge and disk. We fix the density distribution of the bulge and disks by the true values of the simulation (Grand et al., 2017). We have $M_{*,\rm bulge}=3.2\times 10^{10}M_{\odot}$, scale radius $a_{\rm bulge}=1.71$ kpc, and sérsic index $n=1.44$, $M_{*,\rm disk}=5.6\times 10^{10}M_{\odot}$, scale radius $a_{\rm disk}=4.9$ kpc, scale height $b_{\rm disk}=0.25$ kpc for Auriga 23. The disk and bulge mass still have some uncertainty for the MW. We tried models with disk mass $20\%$ higher or lower for Auriga 23 and found no difference in our results. The true values parameterize the bulge and disk of Auriga 5 and Auriga 12 in Grand et al. (2017) are adopted.
We adopt a flexible triaxial generalised NFW model for the DM halo, allowing free orientations of the DM halo, and variable 3D shape as a function of radius generally following Vasiliev et al. (2021):
@@
The coordinate $\boldsymbol{X}\equiv\{X,Y,Z\}$ is related to the galactocentric Cartesian coordinate $\boldsymbol{x}_{\rm gc}\equiv\{x_{\rm gc},y_{\rm gc},z_{\rm gc}\}$ by a rotation matrix parametrised by three Euler angles $\alpha_{q}$, $\beta_{q}$, $\gamma_{q}$ as shown in Section 2.2. In principle, all three angles can be left free. Taking into account the degeneracy between the DM orientation and the 3D shape, we fix $\beta_{q}=0$. The combination of $\alpha_{q}$ and $\gamma_{q}$ determines the DM halo orientation in the $XY$ plane; we fix $\gamma_{q}=0$ and allow $\alpha_{q}$ to be free to test the ability of constraining it. For default models, we also fix $\alpha_{q}=0$.
@@
For the radial profile, we fixed $\alpha=1$, $\beta=3$, chose the outer cutoff radius $r_{\rm cut}=500$ kpc and the cutoff strength $\xi=5$. We start with models for all three mock galaxies with fixed DM orientation ($\alpha_{q}=0$, $\beta_{q}=0$ and $\gamma_{q}=0$) and constant $p_{\rm DM}$ and $q_{\rm DM}$. The five parameters in the DM halo: the scale density $\rho_{0}$ (or the mass $M_{\rm halo}$), the scale radius $r_{s}$, the inner density slope $\gamma$, $p_{\rm DM}$ and $q_{\rm DM}$ are left as free parameters. For Auriga 12, we further construct a model with different orientation angles $\alpha_{q}$ and a model that allows $p(r)$ and $q(r)$ to vary as a function of radius.
@@ 3.2 Orbit integration
Orbit integration is performed using the publicly released AGAMA package (Vasiliev, 2019) [^1] https://github.com/GalacticDynamics-Oxford/Agama. We have about 20,000 stars in the cleaned sample of the smooth halo, with their positions and velocities measured from observations, and the weight of each star obtained from the correction of selection. We start with the position and velocity of the stars from observations and integrate 10 orbital periods for each star. We withdraw 1000 particles from each orbit with equal time steps, and each particle is given the weight as the star that initialises the orbit. We superpose particles drawn from the orbits together, thus obtaining an orbit-superposed model, which numerically represents the DF of the stellar system.
@@ 3.3 Comparing model and data
The DF of the orbit-superposed model, $f_{\rm model}(\boldsymbol{x},\boldsymbol{v}|\rm potential)$, is based on data but incorporates the gravitational potential. If the system is in dynamical equilibrium, the DF of the model with the accurate gravitational potential should be statistically the same as that of the stars that initialised it $f_{\rm data}(\boldsymbol{x},\boldsymbol{v})$. We detail particular components of the DF: the spatial density distribution and velocity distributions at different positions. We evaluated the goodness of the model by comparing these with the data.
We calculate the stellar density distribution in the $R_{\rm gc}$-$z_{\rm gc}$ plane, by dividing it into $25\times 25$ bins within 50 kpc. The density in each bin is calculated by $N/(2\pi RdRdz)$ in units of $N/{\rm kpc}^{3}$. We calculate the density directly for both the observed data $\rho_{\rm data}$ and the model $\rho_{\rm model}$. The uncertainty of the data $d\rho_{\rm data}$ is assessed by bootstrapping the position of stars with their uncertainties. The density distributions of the data and model are shown in Fig. 4; we calculate $\chi^{2}_{\rm den}$ by directly comparing the data and model:
@@
To analyse velocity distributions, we categorise model particles into bins $N_{\rm bin}=7\times 6$ in $r$ and $\theta$. The divisions are $r=[5,8,12,16,22,30,40,50]$ kpc and $\theta$=[0,10,20,30,40,55,90]. We then construct the velocity distributions ($v_{r}$, $v_{\phi}$, $v_{\theta}$) within each bin, as illustrated in Fig. 3. For each velocity component within bin $j$, we represent its distribution by a histogram using intervals $M$, labelled as ($v^{k}_{m,j}$, $h^{k}_{m,j}$), where $k$ corresponds to the velocities $v_{r}$, $v_{\phi}$, and $v_{\theta}$.
@@
The collective likelihood of all $N_{j}$ stars in this bin is given by $L_{j}^{k}=\sum_{i=1}^{N_{\rm j}}\log(P^{k}_{i,j})$, and the overall likelihood across all the bins is represented by $L^{k}=\sum_{j=7}^{N_{\rm bin}}L_{j}^{k}$. Due to a deficit of stars at $r<4$ kpc, which contributes to the distribution function (DF) of the inner bins, the DF for the innermost bins of our model remains incomplete. Therefore, we omit the first six bins within the range r=[5,8] in our likelihood calculations. The likelihood across the three velocity components is merged to derive the constraints from the velocity distributions.
@@ 3.4 Best-fitting models
We investigated the parameter space of $\rho_{0}$, $r_{s}$, $\gamma$, $p_{\rm DM}$, and $q_{\rm DM}$ within the gravitational potential. A parameter grid is established with intervals of 0.1, 5 kpc, 0.02, 0.04, and 0.04 for $\rho_{0}$, $r_{s}$, $\gamma$, $p_{\rm DM}$, and $q_{\rm DM}$, respectively. An iterative approach is used to find the optimal models. The process begins with an initial model, and, after developing initial models, iterative refinement follows. During each iteration, we identify the optimal models using the criterion $\chi^{2}-\min(\chi^{2})<\chi^{2}_{s}$, where $\chi^{2}_{s}=100$. We then generate new models by walking two steps in every direction of the parameter grid of each optimal model. This approach guides the search towards the lower $\chi^{2}$ within the parameter grid and halts when the model with the minimum $\chi^{2}$ is identified. The relatively high value of $\chi^{2}_{s}$ is selected to ensure that all models within $1\sigma$ confidence level are considered before the end of the iteration. Ultimately, we determine the best-fitting models by achieving the minimum $\chi^{2}_{\rm tot}$.
@@
We illustrate how the model works using Auriga 23 as an example. In Fig. 4, it is shown that the best-fitting model accurately reproduces the stellar density across the $R_{\rm gc}-z_{\rm gc}$ plane. Furthermore, it imposes strict constraints on the 3D shape of the underlying DM halo. The shape of the DM halo directly affects the direction of stellar acceleration and thus the shape of the stellar orbits. Variations in the DM halo shape can cause the constructed stellar halo to appear overly flat or excessively round. As shown in Fig. 5, the best-fitting model also reproduces the velocity distributions well in all three components: $v_{r}$, $v_{\phi}$, and $v_{\theta}$. These velocity distributions provide strong constraints on the radial density distribution of the DM halo which directly affects the strength of the acceleration, especially in the radial direction. The different radial profiles of the DM halo lead to notable discrepancies in the velocities between the observational data and the model, particularly in $v_{r}$. Of course, both the stellar density and velocity distributions are determined by the overall 3D DM distribution, which will be better constrained by combining both.
Our best-fitting model generally reproduces both the density and velocity distributions at various locations in the $R_{\rm gc}-z_{\rm gc}$ plane, thus validating our assumption of an overall stationary system with the agreement between the model and the data. However, these stellar system are never in perfect equilibrium. Although we have tried our best to clean the substructures, there will inevitably be some residual substructures or non-equilibrium features left in the data. For example, in Fig. 5, there are some peaks/bumps in the velocity distributions that are not matched by the model, some of them can indicate residual substructures. The $v_{\phi}$ distributions in the bin at $8<r<12$ kpc close to the disk plane are not perfectly matched by the model either, which can indicate some non-equilibrium features.
@@
In Fig. 6, we illustrate the parameter space explored for Auriga 23. The coloured points denote models within a confidence level $3\sigma$, colored by the normalised $\Delta\chi^{2}_{\rm tot}=(\chi^{2}_{\rm tot}-\min(\chi^{2}_{\rm tot}))/\Delta\chi^{2}_{1\sigma}$, where $\Delta\chi^{2}_{1\sigma}$ is the confidence level $1\sigma$ determined by bootstrapping 222 We take the gravitational potential of the best-fitting model, and perturb the position and velocity of observed stars with their uncertainty for 100 trials. In each trial, we recompute the model and evaluate the resulting $\chi^{2}_{\rm tot}$ (including $\chi^{2}_{\rm den}$ and $\chi^{2}_{\rm vdis}$). The standard deviation of these 100 values $\chi^{2}$ is taken as the confidence level $1\sigma$ of the model $\Delta\chi^{2}_{1\sigma}$.. As shown in Fig. 6, the combination of density distribution and velocity distribution strongly constrains the 3D shape of DM $p_{\rm DM}$ and $q_{\rm DM}$. Due to the still limited data coverage, there are still strong degeneracies between the three parameters ($\rho_{0}$, $r_{s}$, and $\gamma$) determining the radial profile. Nevertheless, the radial density profile within the data coverage is well constrained, as will be detailed in the next section.
@@ 4.1 Recovery of DM density profile
We demonstrate how effectively our models recover the DM radial density profile and the enclosed DM mass in Fig. 7. Despite significant degeneracy between $\rho_{0}$, $r_{s}$, and $\gamma$, our models accurately capture the radial density profile within the observational range of 4-50 kpc. The uncertainty is larger at $r<8$ kpc due to the absence of data.
The enclosed DM mass profile is also well recovered, with a relative uncertainty of about $10\%$. It should be noted that the DM mass we obtained in the outer regions is $5\%-10\%$ higher than the true values but aligns well with the enclosed mass that combines DM and gas. The CircumGalactic Medium (CGM) in the halo of the simulation mirrors the DM’s density distribution. The mass of CGM is compensated by an overestimate of the DM mass without a separate gas contribution in our model of the gravitational potential.
@@ 4.2.1 Models with constant $q_{\rm DM}$ and $p_{\rm DM}$
Initially, we developed models with constant $q_{\rm DM}$ and $p_{\rm DM}$ for all three galaxies. As depicted in Fig. 8, our model successfully recovered both $p_{\rm DM}$ and $q_{\rm DM}$ for Auriga 23 and Auriga 5, with relatively small uncertainties. For Auriga 12, the model captured the outer configuration of the DM halo, and there is a significant degeneracy between $p_{\rm DM}$ and $q_{\rm DM}$ when the halo is vertically aligned, characterised by $p_{\rm DM}<q_{\rm DM}$.
The stellar density distribution in the $R_{\rm gc}-z_{\rm gc}$ plane, especially the flattening of the stellar halo, is crucial to constrain the DM shape. We measure the stellar halo flattening, $q_{\rm star}$, across various radii using isodensity contours and compared these from the data and several models in Fig. 9. The stellar halo of Auriga 23 is notably flat, becoming increasingly so from the centre to the outskirts, and our best-fitting model successfully mimicked $q_{\rm star}$ at all radii. The stellar halo of Auriga 5 appears rounder, with a tendency to become rounder moving outward; our model approximates the radial variation $q_{\rm star}(r)$, albeit with moderate inner region discrepancies. The stellar halo of Auriga 12 is the most spherical, notably rounding from inner to outer areas, with $q_{\rm star}(r)$ challenging to replicate by a model with a constant DM shape. The failure to reproduce the tendency of $q_{\rm star}(r)$ suggests a variation of DM shapes from the inner to outer regions.